# Boosting Generative Adversarial Transferability with Self-supervised Vision Transformer Features

### v1

@@ 1 Introduction
Incorporating all three aspects, we introduce dSVA—a generative dual self-supervised ViT features attack. We introduce a novel generative training framework, consisting of a generator to craft transferable adversarial perturbation, and discriminative training approaches to jointly exploit the dual intricate features—both structural and textural—distilled by the two types of self-supervised ViTs. We choose the duo: DINO [7] and MAE [25], for CL and MIM respectively. Figure 1 showcases a birds-eye view of dSVA.
Leveraging the powerful latent representations distilled by self-supervised ViTs, dSVA achieves outstanding adversarial effectiveness. We show an example in Fig. 1 of dSVA successfully disrupting both structural features from DINO (CL) and textural representations from MAE (MIM) (visualized in the attention maps), enabling impressive transferability towards black-box models of distinct architectures. Our experiments demonstrate dSVA’s outstanding transferability to models across ViTs, ConvNets, and MLPs alike, and its ability to evade defenses, surpassing various state-of-the-arts.
@@ 3.2 Facet-level Feature Exploitation
As in Fig. 2, to train $\mathcal{G}_{\bm{\theta}}$ for perturbation generation, dSVA is designed to deviate the latent representations of a benign image and its generated adversarial example, that is, to minimize the cosine similarity between the deep features extracted. In this way, the crafted perturbation would be able to neutralize critical decisive low-level features within the sample, thereby misleading black-box DNNs. Hence, the discriminative loss at this stage is formulated as
@@ 3.3 Self-attention Exploitation
where $H$ is the number of attention heads. Thus, $S^{l}$ serves as a feature landmark guidance for targeting intermediate features, regularizing the global semantic knowledge learnt by the generator. We apply a scaling factor of $\gamma$ to $S^{l}$ for loss optimization. Building on Eq. 2, loss function $\mathcal{L}$ at this stage is thus formulated as
@@
Shown in Fig. 3 is the attention saliency maps extracted from ViTs with self-supervision (red background) vs. supervision (blue background), as well as the variance of saliency maps with increasing layer depth from left to right. Compared to the saliency maps extracted from a supervised ViT, those from the self-supervised ViTs DINO (first row) and MAE (second row) are less noisy and capture various levels of global and local semantics, respectively. From shallow to deep layers, the self-supervised representations favor less spatial information and more textural information, whereas the supervised ViT’s representations collapse into homogeneous primitive patterns. These visualizations showcase the powerful representations offered only by the self-attention of self-supervised ViTs, acting as feature landmarks to be integrated in dSVA for transferability boosts.
@@ 3.4 Joint Self-supervision Feature Discrimination
where $\mathcal{L}_{\text{I}}$ and $\mathcal{L}_{\text{II}}$ are derived as in Eq. 6 from DINO and MAE, respectively. Doing so, dSVA is able to craft highly transferable perturbation that targets both structural and textural image features, greatly boosting transferability across various black-box models with diverse architectures.
@@ 4.1 Experimental Settings
For both DINO and MAE, we choose features extracted at the penultimate layer $l=10$. We select the key facet of DINO and the query facet of MAE to exploit. The joint training parameter of dSVA is set as $\lambda=0.5$. The rationale and empirical evaluations supporting these selections are presented in Secs. 4.4 and 4.5.
@@
Generative attack baselines include BIA [71] and CDA [41]. We use VGG-19 [50], ResNet-152 (Res-152) [23], and DenseNet-169 (Den-169) [27] as their surrogates with the same perturbation budget of $\varepsilon=10$ to follow their setups. We also compare against BIA and CDA trained on supervised ViT-B/16. We additionally include evaluations against gradient-based attacks, including the classic MI-FGSM (MI) [15], and 3 other state-of-the-art attacks designed for ViTs (PNA [62], TGR [70], ATT [38]). (In Tab. 1 and Tab. 2, MI-FGSM is abbreviated as MI so as to avoid confusion with MIM—masked image modeling.)
@@ 4.2 Transferability to Black-box Models
We first evaluate black-box transferability within the ImageNet domain. For attack targets, we choose 3 ConvNets with the same structure as the surrogates of the compared methods to follow baseline settings (VGG-16, ResNet-50 (Res-50), DenseNet-121 (Den-121)). We add 3 ConvNets with a different structure (EfficientNet-B0 (Eff-B0) [55], Inception-v3 (Inc-v3) [53], Inception-v4 (Inc-v4) [54]), 5 ViTs (Swin-B [36], MaxViT-T [58], PiT-B [26], VisFormer-S [11], LeViT-128 [21]), and an MLP Mixer (Mixer-B/16) [56]. We report the results in Tab. 1.
@@ 4.3 Transferability to Defense Models
Next, we validate our approach against defenses, an aspect previously unexplored in the context of generative attacks. We follow previous setups [62, 70, 38] and use 6 robust black-box models on ImageNet to evaluate defense evasion, namely Inc-v3adv, IncRes-v2adv [33], Inc-v3ens3, Inc-v4ens4, IncRes-v2ens [57], and EfficientNet-B0 with AdvProp [65] (Eff-b0ap). Shown in Tab. 2, we once again observe that dSVA shows superior performance across all adversarially trained models, with dSVA (Joint) achieving transferability that exceeds all compared attacks by an average margin of 32.98%. We contend that while adversarial training enhances DNN robustness by developing more resilient features, they ultimately need to use these same essential features for classification. By fully exploiting self-supervised ViT representations, decisive elements of the sample are destroyed at a more generalized level, allowing dSVA to evade these defenses. We provide additional results against state-of-the-art defenses and robust ViTs in Appendix C.
@@ 4.4 Analysis on the Impact of Relevant Parameters
We now turn our focus to dSVA’s deciding parameters, that is, (1) feature facet (query, key, value, or the entire layer’s output: token), (2) feature layer $l$, and (3) $\lambda$, for dSVA (Joint). (In Figs. 4, 5 and 6 of Sec. 4.4, the bold red line represents the mean transferability of the evaluated variant of dSVA, aggregated over observations against target models.)
@@
We first evaluate the performances of dSVA (DINO) and dSVA (MAE) with respect to the exploited facets. We report the black-box transferability of them in Fig. 4. We first note that the variants that directly exploit the token facet, i.e, the entire intermediate layer output, always lags behind, especially in the case of dSVA (MAE). These findings underline the efficacy of our proposed facet-level exploit to capitalize on the adversarial potential of the features distilled by the Transformer architecture. For MAE, the query directly serves as the input with masked patches, which is intuitively more crucial for its reconstruction task. The key facet in this case only provides additional context of the current masked modelling session. This aligns with our observation that dSVA (MAE) performs best with the query facet. For DINO, the student network generates one view of the image as the query, while the teacher uses another as the key. The teacher, acting as a guide, would provide a better contrastive signal. Our results, although not as pronounced as the MAE variant, show that dSVA (DINO) performs best when exploiting the key facet.
@@
Next, we investigate the impact of layer $l$. We report dSVA (DINO) and dSVA (MAE)’s transferability that exploit layer $l$ from 1 to 11 in Fig. 5. We notice that the transferability of dSVA tends to increase as layer deepens. We reason that as the layers deepen, both self-supervised strategies manage to encode richer and more generalizable semantic information, benefiting adversarial transferability. Notably, transferability of both variants drops at the final 11th layer. This is expected as the final layer of Transformer-based models is often optimized for specific training setups, which results in significant reduction in generalizability [14]. In terms of vision tasks, ViTs have also shown to maintain spatial and positional information in all but the last layer [30, 20]. We choose the penultimate layer of $l=10$ of both DINO and MAE in dSVA.
@@
We finally explore the key factor of dSVA (Joint), that is, the balance between feature disruption for DINO (CL) and MAE (MIM), which is controlled by $\lambda$ as described in Eq. 7. The transferability of dSVA (Joint) with $\lambda$ in $(0,1)$ with a step size of 0.1 is reported in Fig. 6. We observe two interesting trends. First, as the dual aspects of features are more incorporated into dSVA (as $\lambda$ approaches the midpoint), adversarial effectiveness increases. This behavior substantiates our hypothesis that the features provided by CL and MIM complement each other under an adversarial context, where both global and local relationships are to be destroyed, highlighting the importance of our proposed joint feature disruption. In addition, as $\lambda$ decreases from 0.9 to 0.5, that is, as the aspect of MIM features increase while CL features decrease, adversarial effectiveness show a tendency to rise. We argue that the while CL provided structures are crucial for shape/object distinction from a human standpoint, to craft generalized perturbation for fooling DNNs, textural details distilled by MIM ought to be more purposefully considered, as DNNs favor these fine-grained details. $\lambda=0.5$ yields the best performances in our setup, but given the similarity of the trends for $\lambda$ in $[0.3,0.5]$, we suggest that the optimal $\lambda$ may vary depending on the specific task or dataset.
@@ 4.5 Visualizing Facet-level Feature Disruption
In Fig. 7, we visualize how self-supervised ViT features are more meaningful than supervised ones, and how some ViT feature facets are more crucial than others. We conduct PCA on DINO, MAE, and supervised ViT-B/16’s features on all facets. We notice once again that the self-supervised features are richer and less noisy than the supervised ones. We find that, for both DINO and MAE, the value and token facets are noisier than the query and key facets. For DINO, its key facet shows more distinct shapes and objects, whereas for MAE, its query facet shows less noisy textured details. These observations align with our parameter selections. We also show how dSVA’s adversarial perturbation equally destroys meaningful semantics within the image, underscoring our approach’s effectiveness in feature disruption.
@@ 4.6 Ablation Study
We finally conduct an ablation study on two factors: (1) self-supervision, and (2) self-attention exploitation. We report the transferability of dSVA with supervised ViT, DINO, MAE, and Joint variants, both w/ and w/o attention saliency map regularization applied, in Fig. 8. For single model variants, we aggregate the results over all facets. For dSVA (Joint), we aggregate the observations over $\lambda\in[0.3,0.5]$.
@@ Appendix A Experimental Details
For dSVA with DINO, MAE, and the vanilla supervised ViT-B/16 at a stride of $s=16$, as well as for all compared generative attacks (CDA, BIA), generator $\mathcal{G}_{\theta}$ is trained on the entirely of the ImageNet training set for one epoch with a batch size of 32. Under this setup, single model variants of dSVA require up to 4 hours of training, a duration comparable to previous methods. For the joint variant, i.e., dSVA (Joint), batch size is set to 22, where its training takes up to 7 hours to complete. Our proposed additional exploit of self-attention (which is optional) in dSVA does not increase the training time. The inference time for the adversarial generator is comparable to, if not faster than, that of gradient-based iterative adversarial attacks. For all settings, GPU memory utilization approximates to over 90%. We organize the rest of the experimental details in Tab. 3, which includes ViTs with stride of $s=8$ that we use in sections that report results of cross-domain transferability.
@@ Appendix B Results of Cross-domain Transferability
Table 4: Transferability towards coarse-grained classification domains. We report transferability (%) towards domains CIFAR-10, CIFAR-100, SVHN, and STL-10. $s$ is the stride of ViT-B/16. $A$ denotes whether attention regularization in dSVA is activated.
Attack $s$ $A$ CUB-200-2011 Stanford Cars FGVC Aircraft
@@
Table 5: Transferability towards fine-grained classification domains. We report transferability (%) towards domains CUB-200-2011, Stanford Cars, and FGVC Aircraft. $s$ is the stride of ViT-B/16. $A$ denotes whether attention regularization in dSVA is activated.
@@
Table 4 showcases our findings on coarse-grained classification domain transferability. With the target models in CIFAR-10 and CIFAR-100 being VGG-like architectures, the BIA attack using a VGG-19 surrogate model unsurprisingly yields superior results. Among the dSVA variants, dSVA (Joint) with DINO and MAE at stride $s=8$ excels, closely matching the baseline performance in these domains. In contrast, for the SVHN and STL-10 domains, dSVA variants outperform the baseline, with dSVA (DINO) surpassing dSVA (Joint) in SVHN due to DINO’s sensitivity to global shape and structure, which aligns with the focus of the SVHN domain on house numbers (digit classification). Interestingly, self-attention exploitation in dSVA does not enhance performance in this coarse-grained context.
Turning to fine-grained classification transferability in Tab. 5, dSVA (Joint) with active self-attention exploitation leads in most scenarios, outperforming nearly all baselines except when the target model is Res-50. Notably, dSVA (DINO) outperforms the otherwise dominant dSVA (Joint) variant in a specific case: attacking the Stanford Cars domain’s SE-Res-101 model.
@@ Appendix C Additional Comparisons of Transferability to Defense Models
In this section, we present additional comparisons on the transferability of dSVA to robust ConvNets, ViTs, and hybrid models with state-of-the-art defenses, which are lacking in prior work. We report the results in Tab. 6, where the citations accompanying the model names refer to the respective state-of-the-art adversarial defenses employed on the model itself. Note that we here use the same experimental setups as in Sec. 4, except for employing a larger $\varepsilon=16$ constraint, otherwise the transferability across all evaluated attacks would be too low to be comparable.

### v2

@@ 1 Introduction
Incorporating all three aspects, we introduce dSVA—a generative dual self-supervised ViT features attack. We introduce a novel generative training framework, consisting of a generator to craft transferable adversarial perturbation, and discriminative training approaches to jointly exploit the dual intricate features—both structural and textural—distilled by the two types of self-supervised ViTs. We choose the duo: DINO [7] and MAE [25], for CL and MIM respectively. Figure˜1 showcases a birds-eye view of dSVA.
Leveraging the powerful latent representations distilled by self-supervised ViTs, dSVA achieves outstanding adversarial effectiveness. We show an example in Fig.˜1 of dSVA successfully disrupting both structural features from DINO (CL) and textural representations from MAE (MIM) (visualized in the attention maps), enabling impressive transferability towards black-box models of distinct architectures. Our experiments demonstrate dSVA’s outstanding transferability to models across ViTs, ConvNets, and MLPs alike, and its ability to evade defenses, surpassing various state-of-the-arts.
@@ 3.2 Facet-level Feature Exploitation
As in Fig.˜2, to train $\mathcal{G}_{\bm{\theta}}$ for perturbation generation, dSVA is designed to deviate the latent representations of a benign image and its generated adversarial example, that is, to minimize the cosine similarity between the deep features extracted. In this way, the crafted perturbation would be able to neutralize critical decisive low-level features within the sample, thereby misleading black-box DNNs. Hence, the discriminative loss at this stage is formulated as
@@ 3.3 Self-attention Exploitation
where $H$ is the number of attention heads. Thus, $S^{l}$ serves as a feature landmark guidance for targeting intermediate features, regularizing the global semantic knowledge learnt by the generator. We apply a scaling factor of $\gamma$ to $S^{l}$ for loss optimization. Building on Eq.˜2, loss function $\mathcal{L}$ at this stage is thus formulated as
@@
Shown in Fig.˜3 is the attention saliency maps extracted from ViTs with self-supervision (red background) vs. supervision (blue background), as well as the variance of saliency maps with increasing layer depth from left to right. Compared to the saliency maps extracted from a supervised ViT, those from the self-supervised ViTs DINO (first row) and MAE (second row) are less noisy and capture various levels of global and local semantics, respectively. From shallow to deep layers, the self-supervised representations favor less spatial information and more textural information, whereas the supervised ViT’s representations collapse into homogeneous primitive patterns. These visualizations showcase the powerful representations offered only by the self-attention of self-supervised ViTs, acting as feature landmarks to be integrated in dSVA for transferability boosts.
@@ 3.4 Joint Self-supervision Feature Discrimination
where $\mathcal{L}_{\text{I}}$ and $\mathcal{L}_{\text{II}}$ are derived as in Eq.˜6 from DINO and MAE, respectively. Doing so, dSVA is able to craft highly transferable perturbation that targets both structural and textural image features, greatly boosting transferability across various black-box models with diverse architectures.
@@ 4.1 Experimental Settings
For both DINO and MAE, we choose features extracted at the penultimate layer $l=10$. We select the key facet of DINO and the query facet of MAE to exploit. The joint training parameter of dSVA is set as $\lambda=0.5$. The rationale and empirical evaluations supporting these selections are presented in Secs.˜4.4 and 4.5.
@@
Generative attack baselines include BIA [71] and CDA [41]. We use VGG-19 [50], ResNet-152 (Res-152) [23], and DenseNet-169 (Den-169) [27] as their surrogates with the same perturbation budget of $\varepsilon=10$ to follow their setups. We also compare against BIA and CDA trained on supervised ViT-B/16. We additionally include evaluations against gradient-based attacks, including the classic MI-FGSM (MI) [15], and 3 other state-of-the-art attacks designed for ViTs (PNA [62], TGR [70], ATT [38]). (In Tab.˜1 and Tab.˜2, MI-FGSM is abbreviated as MI so as to avoid confusion with MIM—masked image modeling.)
@@ 4.2 Transferability to Black-box Models
We first evaluate black-box transferability within the ImageNet domain. For attack targets, we choose 3 ConvNets with the same structure as the surrogates of the compared methods to follow baseline settings (VGG-16, ResNet-50 (Res-50), DenseNet-121 (Den-121)). We add 3 ConvNets with a different structure (EfficientNet-B0 (Eff-B0) [55], Inception-v3 (Inc-v3) [53], Inception-v4 (Inc-v4) [54]), 5 ViTs (Swin-B [36], MaxViT-T [58], PiT-B [26], VisFormer-S [11], LeViT-128 [21]), and an MLP Mixer (Mixer-B/16) [56]. We report the results in Tab.˜1.
@@ 4.3 Transferability to Defense Models
Next, we validate our approach against defenses, an aspect previously unexplored in the context of generative attacks. We follow previous setups [62, 70, 38] and use 6 robust black-box models on ImageNet to evaluate defense evasion, namely Inc-v3adv, IncRes-v2adv [33], Inc-v3ens3, Inc-v4ens4, IncRes-v2ens [57], and EfficientNet-B0 with AdvProp [65] (Eff-b0ap). Shown in Tab.˜2, we once again observe that dSVA shows superior performance across all adversarially trained models, with dSVA (Joint) achieving transferability that exceeds all compared attacks by an average margin of 32.98%. We contend that while adversarial training enhances DNN robustness by developing more resilient features, they ultimately need to use these same essential features for classification. By fully exploiting self-supervised ViT representations, decisive elements of the sample are destroyed at a more generalized level, allowing dSVA to evade these defenses. We provide additional results against state-of-the-art defenses and robust ViTs in Appendix C.
@@ 4.4 Analysis on the Impact of Relevant Parameters
We now turn our focus to dSVA’s deciding parameters, that is, (1) feature facet (query, key, value, or the entire layer’s output: token), (2) feature layer $l$, and (3) $\lambda$, for dSVA (Joint). (In Figs.˜4, 5 and 6 of Sec.˜4.4, the bold red line represents the mean transferability of the evaluated variant of dSVA, aggregated over observations against target models.)
@@
We first evaluate the performances of dSVA (DINO) and dSVA (MAE) with respect to the exploited facets. We report the black-box transferability of them in Fig.˜4. We first note that the variants that directly exploit the token facet, i.e, the entire intermediate layer output, always lags behind, especially in the case of dSVA (MAE). These findings underline the efficacy of our proposed facet-level exploit to capitalize on the adversarial potential of the features distilled by the Transformer architecture. For MAE, the query directly serves as the input with masked patches, which is intuitively more crucial for its reconstruction task. The key facet in this case only provides additional context of the current masked modelling session. This aligns with our observation that dSVA (MAE) performs best with the query facet. For DINO, the student network generates one view of the image as the query, while the teacher uses another as the key. The teacher, acting as a guide, would provide a better contrastive signal. Our results, although not as pronounced as the MAE variant, show that dSVA (DINO) performs best when exploiting the key facet.
@@
Next, we investigate the impact of layer $l$. We report dSVA (DINO) and dSVA (MAE)’s transferability that exploit layer $l$ from 1 to 11 in Fig.˜5. We notice that the transferability of dSVA tends to increase as layer deepens. We reason that as the layers deepen, both self-supervised strategies manage to encode richer and more generalizable semantic information, benefiting adversarial transferability. Notably, transferability of both variants drops at the final 11th layer. This is expected as the final layer of Transformer-based models is often optimized for specific training setups, which results in significant reduction in generalizability [14]. In terms of vision tasks, ViTs have also shown to maintain spatial and positional information in all but the last layer [30, 20]. We choose the penultimate layer of $l=10$ of both DINO and MAE in dSVA.
@@
We finally explore the key factor of dSVA (Joint), that is, the balance between feature disruption for DINO (CL) and MAE (MIM), which is controlled by $\lambda$ as described in Eq.˜7. The transferability of dSVA (Joint) with $\lambda$ in $(0,1)$ with a step size of 0.1 is reported in Fig.˜6. We observe two interesting trends. First, as the dual aspects of features are more incorporated into dSVA (as $\lambda$ approaches the midpoint), adversarial effectiveness increases. This behavior substantiates our hypothesis that the features provided by CL and MIM complement each other under an adversarial context, where both global and local relationships are to be destroyed, highlighting the importance of our proposed joint feature disruption. In addition, as $\lambda$ decreases from 0.9 to 0.5, that is, as the aspect of MIM features increase while CL features decrease, adversarial effectiveness show a tendency to rise. We argue that the while CL provided structures are crucial for shape/object distinction from a human standpoint, to craft generalized perturbation for fooling DNNs, textural details distilled by MIM ought to be more purposefully considered, as DNNs favor these fine-grained details. $\lambda=0.5$ yields the best performances in our setup, but given the similarity of the trends for $\lambda$ in $[0.3,0.5]$, we suggest that the optimal $\lambda$ may vary depending on the specific task or dataset.
@@ 4.5 Visualizing Facet-level Feature Disruption
In Fig.˜7, we visualize how self-supervised ViT features are more meaningful than supervised ones, and how some ViT feature facets are more crucial than others. We conduct PCA on DINO, MAE, and supervised ViT-B/16’s features on all facets. We notice once again that the self-supervised features are richer and less noisy than the supervised ones. We find that, for both DINO and MAE, the value and token facets are noisier than the query and key facets. For DINO, its key facet shows more distinct shapes and objects, whereas for MAE, its query facet shows less noisy textured details. These observations align with our parameter selections. We also show how dSVA’s adversarial perturbation equally destroys meaningful semantics within the image, underscoring our approach’s effectiveness in feature disruption.
@@ 4.6 Ablation Study
We finally conduct an ablation study on two factors: (1) self-supervision, and (2) self-attention exploitation. We report the transferability of dSVA with supervised ViT, DINO, MAE, and Joint variants, both w/ and w/o attention saliency map regularization applied, in Fig.˜8. For single model variants, we aggregate the results over all facets. For dSVA (Joint), we aggregate the observations over $\lambda\in[0.3,0.5]$.
@@ Appendix A Experimental Details
For dSVA with DINO, MAE, and the vanilla supervised ViT-B/16 at a stride of $s=16$, as well as for all compared generative attacks (CDA, BIA), generator $\mathcal{G}_{\theta}$ is trained on the entirely of the ImageNet training set for one epoch with a batch size of 32. Under this setup, single model variants of dSVA require up to 4 hours of training, a duration comparable to previous methods. For the joint variant, i.e., dSVA (Joint), batch size is set to 22, where its training takes up to 7 hours to complete. Our proposed additional exploit of self-attention (which is optional) in dSVA does not increase the training time. The inference time for the adversarial generator is comparable to, if not faster than, that of gradient-based iterative adversarial attacks. For all settings, GPU memory utilization approximates to over 90%. We organize the rest of the experimental details in Tab.˜3, which includes ViTs with stride of $s=8$ that we use in sections that report results of cross-domain transferability.
@@ Appendix B Results of Cross-domain Transferability
Table 4: Transferability towards coarse-grained classification domains. We report transferability (%) towards domains CIFAR-10, CIFAR-100, SVHN, and STL-10. $s$ is the stride of ViT-B/16. AA denotes whether attention regularization in dSVA is activated.
Attack $s$ AA CUB-200-2011 Stanford Cars FGVC Aircraft
@@
Table 5: Transferability towards fine-grained classification domains. We report transferability (%) towards domains CUB-200-2011, Stanford Cars, and FGVC Aircraft. $s$ is the stride of ViT-B/16. AA denotes whether attention regularization in dSVA is activated.
@@
Table˜4 showcases our findings on coarse-grained classification domain transferability. With the target models in CIFAR-10 and CIFAR-100 being VGG-like architectures, the BIA attack using a VGG-19 surrogate model unsurprisingly yields superior results. Among the dSVA variants, dSVA (Joint) with DINO and MAE at stride $s=8$ excels, closely matching the baseline performance in these domains. In contrast, for the SVHN and STL-10 domains, dSVA variants outperform the baseline, with dSVA (DINO) surpassing dSVA (Joint) in SVHN due to DINO’s sensitivity to global shape and structure, which aligns with the focus of the SVHN domain on house numbers (digit classification). Interestingly, self-attention exploitation in dSVA does not enhance performance in this coarse-grained context.
Turning to fine-grained classification transferability in Tab.˜5, dSVA (Joint) with active self-attention exploitation leads in most scenarios, outperforming nearly all baselines except when the target model is Res-50. Notably, dSVA (DINO) outperforms the otherwise dominant dSVA (Joint) variant in a specific case: attacking the Stanford Cars domain’s SE-Res-101 model.
@@ Appendix C Additional Comparisons of Transferability to Defense Models
In this section, we present additional comparisons on the transferability of dSVA to robust ConvNets, ViTs, and hybrid models with state-of-the-art defenses, which are lacking in prior work. We report the results in Tab.˜6, where the citations accompanying the model names refer to the respective state-of-the-art adversarial defenses employed on the model itself. Note that we here use the same experimental setups as in Sec.˜4, except for employing a larger $\varepsilon=16$ constraint, otherwise the transferability across all evaluated attacks would be too low to be comparable.